OpenStudy (anonymous):

Solve triangle ABC if C = 90 degrees A = 7, c = 9, B = 40 degrees

OpenStudy (anonymous):

what to find?

OpenStudy (anonymous):

well you know A = 50 because they have to add up to 180

OpenStudy (lilg132):

a = 50 b = 40 c = 90

OpenStudy (anonymous):

and sin40 = b/9 so b=5.7

OpenStudy (anonymous):

i assume you mean little a (the side) is 7, not the measure of the angle

OpenStudy (anonymous):

since this is a right triangle you can find the measure of the other side using pythagoras. i think this question is not right however

OpenStudy (anonymous):

third side should be \[\sqrt{9^2-7^2}=\sqrt{32}=4\sqrt{2}\]